OpenStudy (anonymous):

One set of traffic lights takes 75 seconds to complete its sequence. Another set takes 120 seconds to complete its sequence. Both sets turn green at precisely 2 pm. How many seconds later do both sets next turn green at the same time?

OpenStudy (anonymous):

Find the least common multiple of 75 and 120, you have how many seconds.

OpenStudy (hoblos):

600

OpenStudy (anonymous):

thank you! whats the easiest way to find the LCM?

OpenStudy (anonymous):

120 and 75 both divide into 600 equally. Therefore 600 seconds = 10 minutes. So 2:10pm.

OpenStudy (anonymous):

\[75=3\times 5^2\] \[120=2^3\times 3\times 5\] \[LCM=2^3\times 3\times 5^2\]

OpenStudy (anonymous):

Yeah I was going to say, prime factorisation is probably the best method. But after a while and lots of practice you tend to just "recognise" numbers and what they go into.

OpenStudy (anonymous):

yup - but prime factorisation is the 'sure fire' way

OpenStudy (anonymous):

do you always ignore the number with a lower power when using prime factorisation, to find out the LCM?

OpenStudy (anonymous):

Here's a couple of useful links on prime factorisation: http://www.helpwithfractions.com/least-common-multiple.html http://www.mathsisfun.com/prime-factorization.html First results from Google ;)

OpenStudy (anonymous):

haha, thank you so much! :P

OpenStudy (anonymous):

if you just want the answer quickly and you are on the computer try this http://www.wolframalpha.com/input/?i=lcm+%2875%2C120%29

OpenStudy (anonymous):

cool link! wish you could get that in exams though..